suppose that the width of a rectangle is three feet shorter than lenght and that the perimeter of the rectangle is 84 feet.
a.)set up an equation for the perimeter involving only L, the leght of the rectangle.
b.)solce the equation algebraically to find the lenght of the rectangle.find the width as well.

Answer by checkley77(12844)

W=L-3
2L+2W=84
2L+2(L-3)=84
2L+2L-6=84
4L=84+6
4L=90
L=90/4
L=22.5 ANS. FOR THE LENGTH.
W=22.5-3=19.5 ANS. FOR THE WIDTH.
PROOF:
2*22.5+2*19.5
45+39=84
84=84